OpenStudy (anonymous):

M(9, 8) is the midpoint of mc136-1.jpg. The coordinates of S are (10, 10). What are the coordinates of R?

OpenStudy (campbell_st):

well the midpoint formula for x is \[x_{m} = \frac{x_{s} + x_{r}}{2}\] so substitute \[x_{m} = 9...and... x_{s} = 10\] so you get \[9 = \frac{10 + x_{r}}{2}\] so to find xr multiply both sides of the equation by 2 and then subtract 10 for y, the formula is \[y_{m} = \frac{x_{s}+x_{r}}{2}\] and y_m = 8 and y_s = 10 solve for y_r hope this helps